# Public Projects with Preferences and Predictions

### v1

@@ 4.2 Prediction Markets
Efficient markets assumption.
Strategic behavior in prediction markets is complex and not fully understood [Chen et al., 2010, Ostrovsky, 2012, Gao et al., 2013, Chen and Waggoner, 2016]. To abstract from the details of strategic behavior, we assume that in the prediction market stage of the combined mechanism, information is fully aggregated.
Definition 3.
We say the efficient markets assumption holds on the combined mechanism if, in any equilibrium, the last participant in the decision market knows and believes $\mathbf{B}$, the true expected external welfare impacts of the alternatives conditioned on all available information.
It is known that in strategic equilibrium of prediction markets, information is always aggregated under a condition called separable securities [Ostrovsky, 2012]. In other words, when information aggregation fails in prediction markets, it is not for strategic reasons, but only because of the expressiveness of the prediction language. Therefore, we believe that efficient markets is a mild assumption on strategic behavior of the agents in the combined mechanism. However, it would not hold when, for example, a participant is the sole holder of information that alternative $k$ has very low welfare, and that participant has a high personal value for choosing $k$.
@@ 4.2.1 Price of Anarchy
We can now analyze Impractical SQUAP with a decision market aggregation mechanism.
@@
Under the efficient market assumption, for any $x>0$, the importance-weighted decision market with the quadratic scoring rule and liquidity parameter $\beta=\epsilon x$ satisfies a deviation bound of $\epsilon^{1/2}x$.
Proof.
Observe that under the efficient market assumption, the decision market is equivalent to the following: run the entire market, where the last participant predicts $\mathbf{B}$. Then, give the last participant one more opportunity to participate; denote their final predictions by $\mathbf{\hat{B}}$. This is equivalent because of the telescoping sum of the market scoring rule, so that the two prediction opportunities have the same net payoff as one opportunity with predicting $\mathbf{\hat{B}}$.
For any participant except the final one, under the efficient market assumption, their strategy does not affect the final predictions $\hat{B}$. One way to see this is by the above equivalence, since the market will end at $\mathbf{B}$ regardless, followed by the last participant’s possible manipulation.
So we only need to analyze the incentives of the final participant $i$. By the decomposition above, we only need to consider the incentives for the final prediction opportunity (since the previous one is fixed to $\mathbf{B}$). The expected utility for $\mathbf{\hat{B}}$ is
$S(\mathbf{\hat{B}};\mathbf{B})-S(\mathbf{B};\mathbf{B})$ $=S(\mathbf{\hat{B}};\mathbf{B})$
$=-\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k})^{2}.$
By the definition of $x$-best-responding in the market, $x>\frac{1}{\beta}\sum_{k}(\hat{B}_{k}-B_{k})^{2}$. So for all $k$, $(\hat{B}_{k}-B_{k})^{2}\leq\beta x=\epsilon x^{2}$, or $|\hat{B}_{k}-B_{k}|\leq\sqrt{\epsilon}x$. ∎
@@
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. Under the efficient markets assumption, Impractical SQUAP with market parameter $\beta=\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
[…]
We observe that, even with constant liquidity $\epsilon=\Theta(1)$, the Price of Anarchy tends to one with the spread. Typically in a prediction market, we have $\epsilon\to 0$ with the size of the market [Abernethy et al., 2014]. We would expect $\epsilon\to 0$ in large settings in practice, but it is not required for good Price of Anarchy. This occurs because the QTM is robust to larger and larger approximations in $\mathbf{W}$ as the spread of the market grows. While one might prefer a larger decision market in practice in hopes that it better aggregates information, this robustness may at least be reassuring. We will also utilize it next to obtain budget balance in many cases.
@@ 4.2.2 Budget balance
Prediction markets are typically assumed to be subsidized. Although the subsidy could come from charging transaction fees in the market, a natural question is whether the transfers of the QTM portion of the mechanism can be used to fund the aggregation mechanism. In general, this is not always achievable, because as shown by Eguia et al. [2019, 2023], the revenue of the QTM stage can shrink to zero even as the spread of the market grows. However, using Proposition 2, we can characterize settings with growing revenue and that therefore allow for a budget balanced mechanism that self-funds its decision market.
@@
[…]
Proof.
We prove that the mechanism can choose liquidity parameters $\beta$ so that the expected revenue is larger than the amount spent on the decision market stage in expectation. In practice, one would like to scale the prediction-market payments by the largest possible amount so that budget balance is achieved. However, we cannot rule out manipulation for reasons similar to the redistribution, so for analysis we assume that the mechanism designer commits to $\beta$ up front. By the telescoping nature of the proper scoring rule, the mechanism’s total spend on the decision market stage in expectation is $S(\mathbf{\hat{B}};\mathbf{B})-S(\mathbf{B^{0}},\mathbf{B})\leq S(\mathbf{B};\mathbf{B})-S(\mathbf{B^{0}};\mathbf{B})=\frac{1}{\beta}\sum_{k}(B_{k}^{0}-B_{k})^{2}$. With at-least-constant disagreement, revenue is at least constant, so $\beta=\Omega(1)$ is possible with ex ante budget-balance. By Corollary 3, in this case and with $T\to\infty$, the Price of Anarchy converges to $1$. ∎
4.3 Wagering Mechanisms
Wagering mechanisms are another common tool for eliciting and aggregating predictions [Lambert et al., 2008, 2015b]. These mechanisms can be preferable to prediction markets because they are one-shot, and thus simpler to implement; moreover, they are budget balanced. However, whereas prediction markets automatically aggregate information, in a wagering mechanism the designer must combine the resulting predictions from a wagering mechanism herself (often by averaging).
We consider a natural extension that does not appear to have been studied: decision wagering mechanisms, analogous to decision markets. Instead of agents wagering on a single future event, we will have them wager on the future event conditional on alternative $k$ being selected, for each $k=1,\dots,m$. Once $k$ is selected and the event observed, we cancel all other wagers and assign payments based on the predictions for $k$.
Specifically, we adapt the Brier betting mechanism [Lambert et al., 2015a] to our setting, defining the importance-weighted quadratic decision wagering mechanism. For each alternative $k$, forecaster $i$ submits an estimate $\hat{b}^{i}_{k}$ for the expected external welfare conditioned on that alternative. Let $\mathbf{\hat{b}}$ consist of all the players’ predictions for all the alternatives. Once an alternative $k\sim\mathbf{p}$ is chosen and the random variable $b^{*}_{k}$ is observed, each player $i$ receives the payoff
$\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k}^{*})=\frac{1}{p_{k}}\left[s(\hat{b}^{i}_{k},b^{*}_{k})-\frac{1}{N}\sum_{j=1}^{N}s(\hat{b}^{j}_{k},b^{*}_{k})\right],$
where $s$ is the quadratic score defined in Equation 9. In other words, forecaster $i$’s payoff consists of her score minus the average score of all participants. Since $\sum_{i}\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k}^{*})=0$, we obtain:
Observation 2.
Wagers and agent loss.
Above, we have considered the special case of a wagering mechanism where all wagers are equal to one. In general, the wagering mechanism weights each participant’s score $s(\hat{b}^{i}_{k},b^{*}_{k})$ by $\frac{y_{i}^{k}}{Y_{k}}$, where $y^{i}_{k}$ is some nonnegative wager on alternative $k$ and $Y_{k}=\sum_{i}y^{i}_{k}$. The wagering mechanism is generally applied in cases where scores $s()$ are bounded in $[0,1]$, so that an agent can be guaranteed to win or lose no more than her wager. In our case, such guarantees can be achieved by adjusting $\beta$, the liquidity parameter of $s$, if an upper bound on $B_{k}^{*}$ is known. Wagering mechanisms can also be subsidized, e.g. by additionally giving each agent a fraction of their score, or by making all payouts nonnegative with a shift. We could use a similar analysis to the prediction market setting to use the QTM revenue to subsidize the wagering mechanism.
Strategic behavior and aggregation.
We adopt the standard wagering mechanism model of immutable beliefs, where each agent $i$ has fixed beliefs $\mathbf{b}^{i}$ about the expectation of $\mathbf{b}^{*}$, regardless of others’ beliefs. (In a Bayesian model, we essentially have a “no-trade” situation in which the zero-sum payoffs discourage participation.) The wagering mechanism in isolation is strategyproof: agents maximize expected net payoff by reporting their true believed estimates. This follows immediately from the properness of the scoring rule $s$ and the fact that in $\pi_{i}$, the agent can only affect their own score. For analysis, we suppose that the wagering mechanism outputs $\hat{B}_{1},\dots,\hat{B}_{m}$ where $\hat{B}_{k}=\frac{1}{N}\sum_{i=1}^{N}\hat{b}_{k}^{i}$. We assume that, if all agents are truthful, then the output $\mathbf{B}$ is the true expectation of the external welfare impacts. Our analysis could be adapted to other aggregation methods than taking the average.
4.3.1 pPoA guarantee
It only remains to show the deviation bound. Because the wagering mechanism relies on the same quadratic scoring rule as the prediction market, the analysis (Appendix B) uses the same tools, but is simpler.
Lemma 6.
The importance-weighted quadratic decision wagering mechanism with parameter $\beta=\epsilon x$ has a deviation bound $\epsilon^{1/2}x$.
Corollary 5.
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. Impractical SQUAP with the importance-weighted quadratic decision mechanism with wagering parameter $\beta=\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
[…]
@@ 5 Discussion
This paper’s first contribution was a Price of Anarchy analysis of the Quadratic Transfers Mechanism, primarily in the two-alternative setting. Prior work had welfare analysis in asymptotic settings with randomly drawn agents. But results for such a setting, reminiscent of a “strategic jury theorem”, left open a question of robust guarantees with arbitrary sets of agents, as we gave in Section 2.
The second contribution was to extend the QTM to a case with external welfare impacts of the decisions. Many public projects settings involve impacts on non-decisionmakers, such as the climate impacts of a particular policy. Section 3 proposed a practical mechanism for this problem, the Synthetic Players QTM. We were able to show strong Price of Anarchy guarantees for a less-practical variant in which the decisionmaker needs to estimate the total values of each alternative. It approximately maximizes the welfare of decisionmakers plus externalities of the decision.
The third and main contribution was to propose decisionmaking mechanisms that combine preference aggregation with information aggregation. We model the information aggregation stage as forecasting of the external welfare impacts of the decision. By using either prediction markets or wagering mechanisms, we again showed that a less-practical variant of the Synthetic Players QTM with Predictions (SQUAP) has strong Price of Anarchy guarantees. The importance-weighting technique gave forecasters the same expected payoff for any decision of the mechanism. The proper scoring rules enforced penalties for inaccurate predictions, which can only change the final decision with a small probability while causing a prohibitive penalty.
@@
There are a number of technical open problems. First is improving the QTM Price of Anarchy guarantee for more than two alternatives, or giving counterexamples for parameter regimes where it is impossible. Second is proving a Price of Anarchy guarantee for the Synthetic Players QTM (i.e. the practical variant), or finding a counterexample. Even in the latter case, we expect a “Price of Stability” result to be possible. Third is to extend SQUAP to other information-aggregation mechanisms, perhaps with a more practical method than importance weighting, and prove guarantees for the practical variant.
@@ A.1 Equilibrium Characterization
This section contains proofs and discussion from Section 2.2
@@
Fix all other agents’ strategies as $\mathbf{a}^{-i}$. Denote the expected utility $u^{i}$ of pure strategy $\mathbf{a}^{i}$ given fixed strategies $\mathbf{a}^{-i}$ as $u^{i}(\mathbf{a}^{i};\mathbf{a}^{-i})$, and the softmax probability $p_{k}$ over strategy $a^{i}$ and fixed strategies $\mathbf{a}^{-i}$ as $p_{k}(\mathbf{a}^{i};\mathbf{a}^{-i})$. Consider the strategy $\hat{a}_{k}^{i}=0\;\forall k\in\mathbb{Y}$. Then
$u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}0^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (13)
@@
Now, take a different pure strategy $\mathbf{\tilde{a}}^{i}$ for voter $i$ where $|\tilde{a}_{k}^{i}|>\sqrt{\frac{max_{l}v_{l}^{i}}{c}}$ for some $k\in\mathbb{Y}$. Since $(\tilde{a}_{k}^{i})^{2}$ is nonnegative, we must have $\sum_{k=1}^{m}(\tilde{a}_{k}^{i})^{2}>\frac{1}{c}(\max_{l}v_{l}^{i})$. Thus
$u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}(\mathbf{\tilde{a}}_{k}^{i})^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (19)
$<\max_{l}v_{l}^{i}-\max_{l}v_{l}^{i}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$ (22)
@@
Combining Inequalities 16 and 25, for any strategy profile $\mathbf{a}^{-i}$ of voters other than $i$, $u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})<\frac{1}{n-1}\sum_{j\neq i}\sum_{k=1}^{m}(a_{k}^{j})^{2}<u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$. By definition, then, it is a strictly dominated strategy for any agent $i$ to play $\mathbf{\tilde{a}}^{i}$. ∎
@@
Recall that Lemma 1 stated that if $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then agent utilities are strictly concave on $\operatorname{\mathbb{R}}^{m}$.
Proof of Lemma 1.

### v2

@@ 5.4.1 Price of Anarchy
We now consider the welfare guarantees of SQUAP with the decision market aggregation mechanism. We leave the proof, which follows a similar pattern as the analogous result for wagering mechanisms, in Appendix C.
@@
Under the efficient market assumption, for any $x>0$, the importance-weighted decision market with the quadratic scoring rule and liquidity parameter $\beta=\epsilon x$ is $(\epsilon^{1/2},x)$-robust.
@@
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. In the two alternative setting and under the efficient markets assumption, SQUAP with market parameter $\beta=\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
$\text{pPoA}(\mathbf{v},\mathcal{D})\geq 1-\frac{2\epsilon^{1/2}}{T}-\left(\frac{4}{T}\right)^{2/5}.$
As in the wagering mechanism setting, even with constant liquidity $\epsilon=\Theta(1)$, the Price of Anarchy tends to one with the spread. Typically in a prediction market, we have $\epsilon\to 0$ with the size of the market (Abernethy et al., 2014). We would expect $\epsilon\to 0$ in large settings in practice, but it is not required for good Price of Anarchy. This occurs because the QTM is robust to rougher estimates as the spread of the market grows. While one might prefer a larger decision market in practice in hopes that it better aggregates information, this robustness may at least be reassuring. We will also utilize it next to obtain budget balance in many cases.
@@ 5.4.2 Budget balance
Prediction markets are typically assumed to be subsidized. Although the subsidy could come from charging transaction fees in the market, a natural question is whether the transfers of the QTM portion of the mechanism can be used to fund the aggregation mechanism. In general, this is not always achievable, because as shown by Eguia et al. (2019, 2023), the revenue of the QTM stage can shrink to zero even as the spread of the market grows. However, Proposition 7 allows us to characterize settings with growing revenue, therefore leading to a budget balanced mechanism that self-funds its decision market.
@@
Let the disagreement of a value profile $\mathbf{v}$ be $D:=\frac{\sum_{i}(v_{1}^{i}-v_{2}^{i})^{2}}{(W_{1}-W_{2})^{2}}$. Fixing the information structure $\mathcal{D}$, in a sequence of SQUAP mechanisms with growing spread $T\to\infty$ and bounded-below disagreement $D\geq\Omega(1)$, the mechanism does not lose money in expectation and $\text{pPoA}\to 1$.
@@ 6 Discussion
This paper’s first contribution was a Price of Anarchy analysis of the Quadratic Transfers Mechanism, primarily in the two-alternative setting. Prior work analyzed welfare in asymptotic settings with randomly drawn agents. But results for such a setting, reminiscent of a “strategic jury theorem,” left open a question of robust guarantees with arbitrary sets of agents, as we gave in Section 3.
The second contribution was to extend the QTM to a case with external welfare impacts of the decisions. Many public projects settings involve impacts on non-decisionmakers, such as the climate impacts of a particular policy. Section 4 proposed a mechanism for this problem, the Synthetic Players QTM, for which we provided strong Price of Anarchy guarantees. In particular, the mechanism approximately maximizes the welfare of decisionmakers plus externalities of the decision. We also suggest two variants of the mechanism which could be more practically implemented by the designer.
The third and main contribution was to propose decisionmaking mechanisms that combine preference aggregation with information aggregation. We model the information aggregation stage as forecasting of the external welfare impacts of the decision. By using either prediction markets or wagering mechanisms, we again showed that the the Synthetic Players QTM with Predictions (SQUAP) has strong Price of Anarchy guarantees. The importance-weighting technique gave forecasters the same expected payoff for any decision of the mechanism. The proper scoring rules enforced penalties for inaccurate predictions, which can only change the final decision with a small probability while causing a prohibitive penalty.
@@
There are a number of technical open problems. First is improving the QTM Price of Anarchy guarantee for more than two alternatives, or giving counterexamples for parameter regimes where it is impossible. Second is proving a Price of Anarchy guarantee for the practical variants of the Synthetic Players QTM, or finding counterexamples. Even in the latter case, we expect a “Price of Stability” result to be possible. Third is to extend SQUAP to other information-aggregation mechanisms, perhaps with a more practical method than importance weighting, and prove guarantees for the variant.
@@ A.1 Equilibrium Characterization
In this section, we characterize conditions under which a pure-strategy equilibrium exists in the QTM. Our results build on some techniques of  Eguia et al. (2019), but also introduce new techniques and new structural features of the mechanism, such as the fact that votes across alternatives always sum to zero in equilibrium.
To show existence, we take the following steps. First (Lemma 5), we show that for each player, it is strictly dominated to play any $\mathbf{a}^{i}$ lying outside a certain compact set. It follows that all equilibria are “as if” the game were restricted to a compact strategy space (Lemma 6). From there, we can show that if the parameter $c$ is large enough relative to the maximum agent preference, then utilities are concave, and apply an existence theorem (Proposition 4). A similar argument appears in Eguia et al. (2019), but in our case we give a constructive bound on $c$ that appears quite tight. We also note a connection in the analysis to exponential family distributions. Finally, we derive the first-order conditions for pure-strategy Nash equilibrium, and show as a byproduct that each agent’s votes sum to zero.
@@
Fix all other agents’ strategies as $\mathbf{a}^{-i}$. Denote the expected utility $u^{i}$ of pure strategy $\mathbf{a}^{i}$ given fixed strategies $\mathbf{a}^{-i}$ as $u^{i}(\mathbf{a}^{i};\mathbf{a}^{-i})$, and the softmax probability $p_{k}$ over strategy $a^{i}$ and fixed strategies $\mathbf{a}^{-i}$ as $p_{k}(\mathbf{a}^{i};\mathbf{a}^{-i})$.
Consider the strategy $\hat{a}_{k}^{i}=0\;\forall k\in\mathbb{Y}$. Then
$u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}0^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$
@@
Now, take a different pure strategy $\mathbf{\tilde{a}}^{i}$ for voter $i$ where $|\tilde{a}_{k}^{i}|>\sqrt{\frac{max_{l}v_{l}^{i}}{c}}$ for some $k\in\mathbb{Y}$.
Since $(\tilde{a}_{k}^{i})^{2}$ is nonnegative, we must have $\sum_{k=1}^{m}(\tilde{a}_{k}^{i})^{2}>\frac{1}{c}(\max_{l}v_{l}^{i})$. Thus
$u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})$ $=\sum_{k=1}^{m}p_{k}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})v_{k}^{i}-c\sum_{k=1}^{m}(\mathbf{\tilde{a}}_{k}^{i})^{2}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$
$<\max_{l}v_{l}^{i}-\max_{l}v_{l}^{i}+\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{m}(a_{k}^{j})^{2}$
@@
Combining Inequalities 8 and 9, for any strategy profile $\mathbf{a}^{-i}$ of voters other than $i$, $u^{i}(\mathbf{\tilde{a}}^{i};\mathbf{a}^{-i})<\frac{1}{n-1}\sum_{j\neq i}\sum_{k=1}^{m}(a_{k}^{j})^{2}<u^{i}(\mathbf{\hat{a}}^{i};\mathbf{a}^{-i})$. By definition, then, it is a strictly dominated strategy for any agent $i$ to play $\mathbf{\tilde{a}}^{i}$.
@@
Lemma 7.
Lemma 7.
@@
Now, we have the tools to prove Proposition 4, that for $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ a pure-strategy Nash equilibrium exists:
Proposition 4.
Proposition 4.
By Lemma 7, if the mechanism chooses $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ then agent utilities are strictly concave as a function of their votes. Since the strategy spaces of each agent are non-empty, compact, convex subsets of a Euclidean space; utilities are continuous; and utilities are strictly concave, it follows that a pure strategy Nash equilibrium exists (Glicksberg, 1952; Debreu, 1952). ∎
@@
We now state the important first-order conditions that characterize equilibria in terms of the mechanism’s expected individual and aggregate welfare, $\operatorname{\mathbb{E}}_{k\sim p}v_{k}^{i}=\sum_{k}p_{k}v_{k}^{i}$ and $\operatorname{\mathbb{E}}_{k\sim p}V_{k}=\sum_{k}p_{k}V_{k}$ respectively. This result also appears throughout Eguia et al. (2019), although we add some observations that are useful to us later in constructing equilibria as part of our mechanisms.
Lemma 8.
$A_{k}$ $=\frac{p_{k}}{2c}\left(V_{k}-\sum_{\ell}p_{\ell}V_{\ell}\right).$ (13)
Furthermore, if the mechanism sets $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then $\mathbf{a}$ is a pure-strategy Nash equilibrium if and only if (12) and (13) are satisfied.
Furthermore, for any vote totals $\mathbf{V}$, if $\mathbf{A}$ is any aggregate vote vector that satisfies (13), then for any $\mathbf{v}$ summing to $\mathbf{V}$ there exists a corresponding $\mathbf{a}$ summing to $\mathbf{A}$ satisfying (12).
Summing over Equations 12 and 13, we obtain:
Corollary 6.
An equilibrium is unique if Equation 13 has a unique solution $\vec{A}$. In particular, we now prove Proposition 5, which states that equilibria are unique when in the two alternative setting:
Proposition 5.
In the QTM on agent value profile $\mathbf{v}$ for $m=2$ alternatives, if the mechanism chooses $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$, then there is a unique pure-strategy Nash equilibrium.
Proposition 5.
Fix all values $\mathbf{v}$ and set $c\geq\frac{1}{2}\max_{i,k}v_{k}^{i}$. We proceed by contradiction: assume two distinct equilibrium solutions exist $(p_{1},p_{2})$ and $(\hat{p}_{1},\hat{p}_{2})$. WLOG, assume $p_{1}>\hat{p}_{1}$. Since $V_{1}>V_{2}$, by Corollary 7 we know $p_{1}>\hat{p}_{1}>1/2$. It follows that $A_{1}>\hat{A}_{1}$ since the function $\frac{1}{1+e^{-2x}}$ is monotone increasing and $p_{k}=\frac{1}{1+e^{-2A_{k}}}$ for $k=1,2$.
By FOC (Corollary 7), then, $A_{1}-\hat{A}_{1}=(p_{1}p_{2}-\hat{p}_{1}\hat{p}_{2})\frac{1}{2c}(V_{1}-V_{2})>0$. Since $V_{1}-V_{2}>0$, it follows that $p_{1}p_{2}-\hat{p}_{1}\hat{p}_{2}>0$. In other words, $p_{1}(1-p_{1})>\hat{p}_{1}(1-\hat{p}_{1}).$ However, the function $x(1-x)$ is monotone decreasing for $x\geq 1/2$, so this is a contradiction since we assumed $p_{1}>\hat{p}_{1}>1/2$. The result follows. ∎